NADP/AIRMoN Report


Sample # AA4764L is in the refrigerator.

Samples through # AA4748L have been analyzed for pH and conductivity
Samples through #AA4748L have been analyzed by FIA for NH4 and PO4
Samples through #AA4643L have been analyzed by IC for Cl NO3 and SO4
Samples through # AA4626L have been analyzed by AA for Ca, Mg, Na, and K
Samples through #AA4521L have been sent to the sites as preliminary data reports. This
means samples through:

8/6 for DE02
8/9 for IL11
8/10 for MD15
8/8 for NY67
8/6 for OH09
8/4 for PA15
8/6 for TN00
8/6 for VT99

The next preliminary data report should be going out the first of Nov. with data through
#AA4597L or samples through:

DE02 9/3
FL18 9/2
IL11 9/3
MD15 8/22
NY67 8/27
OH09 9/5
PA15 9/5
TN00 9/4
VT99 9/3

We started receiving samples from FL18 with the first sample date off being: 8/12/96.

VT99 has been refunded and is now back in operation, as of the first of June. Down time approximately April and May.

MD15 was flooded during the hurricane season and the sample (and some of the equipment) was lost. Repairs to the equipment have been made and the site is back on line.

MD15 operator, Hoss Parks, OH09 operator, Nicole Bolton, and VT99 operator, Joanne Cummings all attended the NADP/NTN Training Course and got to see the new AlRMoN E-FOF demonstrated.

The e-FOF is expected to be on line at IL11, OH09, and DE02 by the end of Oct. It has been tested by Van, Lori, and myself here at the CAL and is now waiting for the site operators to field test it. If all goes well, it should be on line at 8 of the 9 AIRMoN sites by the first of the year. MD15, due to the isolation of Smith Island, does not have e-mail capabilities so will continue to use the FOF as it is now.

So far we have reported 4205 samples to CSU.
2775 samples are Wet samples (W)
222 samples are Wet but with insufficient sample for complete analysis (WI)
745 samples or bucket changes with no precip. or insufficient sample to send to the
CAL (D)
196 field blanks poured into the bucket (DF)
194 field blanks not poured into the bucket (DK)
73 internal blind samples

Samples through #AA4202L have been sent to CSU with dates through:

6/11 for DE02
6/14 for IL11
6/5 for MD15
6/10 for NY67
6/12 for OH09
6/13 for PA15
6/12 for TN00
6/11 for VT99

The next set of data for CSU has begun to be processed and will include data through the middle of July. It is expected that data through the middle of Aug. will be at CSU by the end of Nov., giving us a 3 month turn around for AIRMoN data.

Three more data reports should be available at the CO by the end of the year covering data through September, 1996.

Samples from NY67 through December, 1993 have been discarded. We are waiting to hear from the other sites as to whether we can discard their samples through at least 1993 as we have no room in the refrigerator. Samples for 1994 will be discarded in 1997, as per the agreed upon protocols.

The sites will be winterizing their raingages on or before Oct. 22.

We are averaging about 110 samples per month although that load should increase to about 130 samples per month now that all 9 sites are operational.